OpenStudy (bealiberty47):

Write a function for the following and please explain how to figure it out. X|2|4|6 Y|1|0|-1

OpenStudy (bealiberty47):

@Kainui @mathmate @amistre64

OpenStudy (bealiberty47):

@xapproachesinfinity

OpenStudy (amistre64):

thats already written as a function

OpenStudy (amistre64):

\[f=\begin{pmatrix}2&4&6\\ 1&0&-1 \end{pmatrix} \]

OpenStudy (bealiberty47):

Sorry I meant a function rule

OpenStudy (amistre64):

how about we make some switches, then we can define whatever we want for these points

OpenStudy (bealiberty47):

ok how

OpenStudy (amistre64):

well s1 = (1 + (x-1.5)/|x-1.5|)/2 is 0 if x < 1.5 and 1 if x > 1.5 s2 = (1 + (x-3.5)/|x-3.5|)/2 is 0 if x < 3.5 and 1 if x > 3.5 s3 = (1 + (x-5.5)/|x-5.5|)/2 is 0 if x < 5.5 and 1 if x > 5.5 f(x) = 1*s1 -1*s2 -1*s3 should work fine

OpenStudy (amistre64):

y = (1 + (x-1.5)/|x-1.5|)/2 - (1 + (x-3.5)/|x-3.5|)/2 - (1 + (x-5.5)/|x-5.5|)/2 http://www.wolframalpha.com/input/?i=y+%3D+%281+%2B+%28x-1.5%29%2F%7Cx-1.5%7C%29%2F2+-+%281+%2B+%28x-3.5%29%2F%7Cx-3.5%7C%29%2F2+-+%281+%2B+%28x-5.5%29%2F%7Cx-5.5%7C%29%2F2

OpenStudy (amistre64):

in other words from -infinity to 1.5, the s1, s2, s3 = 0 from 1.5 to 3.5, the s1 = 1 from 3.5 to 5.5, the s2 turns on and we get (1-1) = 0 from 5.5 to infinity, the s3 turns on and we get (1-1-1) = -1

OpenStudy (amistre64):

or, you can be boring and just make a line ... your choice

OpenStudy (amistre64):

now you tell me, which one is more impressive to figure out :)

OpenStudy (bealiberty47):

Brah this is 8th grade

OpenStudy (bealiberty47):

I need you to put this in laments terms @amistre64

OpenStudy (bealiberty47):

ok nevermind I saw the attachment I get it

OpenStudy (amistre64):

both are functions that satisfy the points .... ones just more blah than the other lol

OpenStudy (bealiberty47):

Thanks for explaining it to me :)